OpenStudy (anonymous):

find the limits if they exist and find the indicated value. Let y(x)=|3x|/x a) lim y(x) as x approaches 0+ B) lim y(x) as x approaches 0- C)lim y(x) as x approaches 0 D) y(0) pleas show step bye step please n explain well.. thanks =)

OpenStudy (anonymous):

A)3 B)-3 C)Undefined D)Undefined

OpenStudy (anonymous):

A)Evaluate for x = 1 B)x = -1 C)Since A does not = B, dose not exist D)Division by zero error.

myininaya (myininaya):

\[x>0 =>|3x|=3x, x<0=>|3x|=-3x\]

OpenStudy (anonymous):

function is \[f(x) = \frac{|3x|}{x} = \left\{\begin{array}{rcc} 3& \text{if} & x > 0 \\ -3& \text{if} & x < 0 \end{array} \right. \] so as you go to 0 from the left you get -3 from the right you get 3 and since those numbers are not the same the limit does not exist at 0

OpenStudy (zarkon):

that is a good looking piecewise function :)

myininaya (myininaya):

yep

OpenStudy (anonymous):

those be the steps, and props to zarkon for letting me still his latex

OpenStudy (anonymous):

*steal

myininaya (myininaya):

i like i didn't know you could do that with latex

OpenStudy (anonymous):

i actually found it on the web but zarkon added the nice text box

OpenStudy (anonymous):

how was work?

myininaya (myininaya):

i'm sleepy since i had to wake up early but work was okay (boring meeting) i come back later and play with it

myininaya (myininaya):

watching movie :)

OpenStudy (anonymous):

good movie i hope. enjoy

OpenStudy (anonymous):

First of all remember the definition of absolute value: \[\left| x \right| = x \] for \[x > 0\] and \[\left| x \right| = -x\] for \[x < 0\] therefore you have that \[\left| 3x \right|/x = 3x/x = 3\] for \[x > 0\] and \[\left| 3x \right|/x = -(3x)/x = -3\] for \[x <0\] Therefore the lim as x tends to \[0^{+}\] we just look which x satisfies that and clearly is the one which has \[x >0\] therefore it is 3. Now the lim as x tendo to \[0^{-}\] we just look which x satisfies that and clearly is the one which has \[x < 0\] which is -3. Therefore since both limits are different then the lim as x tend to 0 is not defined.

OpenStudy (zarkon):

little slow there jobh8810....but I'll give you a medal anyway ;)

OpenStudy (anonymous):

sorry I'm new at this so I'm still learning how to use it hahaha, Peace, Love and Happiness from Puerto Rico

OpenStudy (anonymous):

thanks

OpenStudy (zarkon):

Peace, Love and Happiness from Pennsylvania :)

OpenStudy (anonymous):

nice, I was at Penn State University last August for the MASS program, it was pretty tought but I loved it

OpenStudy (zarkon):

nice

OpenStudy (anonymous):

from pa? nice

OpenStudy (zarkon):

not originally. I moved to PA for my current position.

OpenStudy (anonymous):

why r u still looking at this?

OpenStudy (anonymous):

myin could you help me out real quick?